OpenStudy (anonymous):

"Find the perimeter of the right triangle." one of the legs are 12in, the other is 9in. I dont understand how im supposed to find the hypotenuse to get the perimeter. please help?

jigglypuff314 (jigglypuff314):

pythagorean theorem (hypotenuse)^2 = (one side)^2 + (other side)^2 so hypotenuse^2 = 12^2 + 9^2 so hpyotenuse = √(144 + 81) = √(225) = 15

OpenStudy (amistre64):

try to avoid doing all the work like that. let them get their hands dirty, then correct them if they make a mistake, or encourage them when they get it right :)

jigglypuff314 (jigglypuff314):

doesn't really work if they're afk...